OpenStudy (anonymous):

Help Please 4x/9 + x-10/9

OpenStudy (anonymous):

what is your question

OpenStudy (anonymous):

I have to Simplify it

OpenStudy (anonymous):

please write the question in a non confusing way, where does the division end or start, please put things in bracket :)

OpenStudy (anonymous):

(4x/9) + (x-10/9) then "/" = divide

OpenStudy (anonymous):



OpenStudy (anonymous):

ok thats looking good, sorry I was also helping others,

OpenStudy (anonymous):

ok, because both terms have the same denominator then just plus the 4x and x together

OpenStudy (anonymous):

4x^2?

OpenStudy (anonymous):

x+x=2x x times x= x^2 hope this hints to you

OpenStudy (anonymous):

Ok im confused now.

OpenStudy (anonymous):

so 4x+x would equal 2x?

OpenStudy (anonymous):

ok, 4x+1x= 1x=x

OpenStudy (anonymous):

so it equals 5x

OpenStudy (anonymous):

ohh okay! so now, 5x-10 would be -5x over 9?

OpenStudy (anonymous):

 this is not the end of it